OpenStudy (s3a):

Integrating sin(x) * sin(x+1) dx I think we need to use integration by parts since this problem is under the integration by parts section. Could someone please explain to me how to integrate this problem because, when I do it (using integration by parts), I get a meaningless tautology where sin(1) = some_constant!

OpenStudy (s3a):

You don't need to do all the work; just telling me what to do, in a few words, would be good enough.

OpenStudy (amistre64):

sin(a+b) = sina cosb + sinb cosa might be useful ... but might not :)

OpenStudy (s3a):

Actually, I just found sin(a) sin(b) = 1/2 cos(a-b) - 1/2 cos(a+b), which worked well for me. Thanks, nonetheless! :)

OpenStudy (amistre64):

thatll work fin too :)

OpenStudy (s3a):

Alright, thanks again. See you later.

OpenStudy (anonymous):

\[2\sin a \sin b=\cos \left( a-b \right)-\cos \left( a+b \right)\]